OpenStudy (knudsona):

A box has dimensions of 17 inches long, 1.3 feet wide, and 8 inches high. What is the volume of the box? The formula for the volume is V = l • w • h.

OpenStudy (whpalmer4):

Okay, you need to decide whether you want to do the computation in cubic inches or cubic feet. As you have a mixture of units, you need to convert so that you are using consistent units (all inches, all feet, all meters, whatever). If you want to convert a measurement in inches to one in feet, multiply the quantity in inches like this: \[3 \text{ feet }*\frac{12 \text{ inches }}{1 \text{ foot }} = 3\cancel{\text{ feet }}*\frac{12 \text{ inches }}{1 \cancel{\text{ foot }}} = 36 \text{ inches }\]

OpenStudy (anonymous):

V=26 inches i think

OpenStudy (whpalmer4):

Going from inches to feet: \[36 \text{ in } * \frac{1 \text{ft}}{12 \text{ in }} = 3 \text{ ft} \]

OpenStudy (whpalmer4):

@theonlycool13 No, that is not correct. "inches" is not a unit of volume measurement.